Factor each expression
a2(b - c)- 16b2 (b - c)

Thank You!!

Answer by fractalier(6550)

Okay, factor out the (b-c) first, then factor what remains...so we have
a^2(b - c) - 16b^2 (b - c) =
(b - c)(a^2 - 16b^2) =
(b - c)(a + 4b)(a - 4b)